# Your Classifier Can Do More: Towards Bridging the Gaps in Classification, Robustness, and Generation

### v1

@@ 4.3 Ablation Study
Table 5: Ablation study of EB-JDAT-JEM++, ECO denotes the epoch at which collapse occurs.
@@
1 0 0.1 0.9 90.61 66.41 64.71 146.02 4.44 n/a
2 0.5 0 0.5 74.16 46.28 43.56 69.37 5.95 20
3 0.9 0.1 0 12.59 11.59 11.48 74.29 5.17 44
4 0.35 0.25 0.35 87.13 60.31 58.20 48.52 6.56 70
5 0.45 0.05 0.45 91.43 65.69 64.34 38.89 7.66 n/a
In EB-JDAT, we employ a weighted combination of gradients, as detailed in Algorithm 1, to compute the overall optimization gradient, expressed as $h_{\theta}=w_{1}h_{1}+w_{2}h_{2}+w_{3}h_{3}$. To assess the individual contribution of each weight, we perform ablation studies by setting each of $w_{1}$, $w_{2}$, and $w_{3}$ to zero and evaluating their impact on overall performance. The results, presented in Table 5, reveal two key findings: (1) When $w_{2}=0$, the optimization focuses solely on classification loss and the clean-sample distribution. Although this setting enhances robustness, the significant energy gap between clean and adversarial samples prevents model convergence, leading to model collapse by the 20th epoch. (2) Maintaining a small value for $w_{2}$ not only balances accuracy and robustness but also stabilizes the training phase, reducing the risk of model collapse. Consequently, experimental results indicate that $w_{1}=0.45$, $w_{2}=0.05$, and $w_{3}=0.45$ yields optimal performance in terms of accuracy, robustness, generative capability, and stability. This configuration is therefore adopted as the default setting.
5 Limitation And Broader Impact
Training EB-JDAT on complex, high-dimensional data remains challenging. This challenge is also encountered by JEMs,including JEM grathwohl2019your , JEM++ yang2021jem++ and SADAJEM yang2023towards . This instability arises from the typically sharp probability distribution of real data in high-dimensional space, which leads to inaccurate guidance for image sampling in regions with low data density. Although the approach we propose is a general and flexible optimization framework for all JEMs, considering training stability, we recommend training within faster and more stable JEM variants, such as JEM++ and SADAJEM. Nevertheless, our approach allows a standard classifier to achieve good generative ability and robustness without sacrificing much accuracy, resulting in the best overall performance in terms of accuracy, generative capability, and robustness.
From a societal perspective, a robustness generative-classifier more effectively mitigates data scarcity and adversarial perturbations. However, if the training dataset itself is high-risk, the model may generate equally high-risk content, potentially posing harm to society. To minimize such harm, we will soon release user guidelines and begin implementing safety filters.
@@ 6 Conclusion
In this work, we systematically investigate the underlying causes of the performance gap between discriminate classifiers, robust classifiersmkadry2017towards ; zhang2019theoretically and JEM-based classifiers grathwohl2019your ; yang2023towards in terms of accuracy, generative capability, and adversarial robustness, and propose a general and flexible optimization framework, called EB-JDAT to eliminate the inherent trilemma. Our experiments validate the effectiveness of these techniques across various benchmarks, demonstrating superior performance in achieving the triple trade-off, with particularly impressive SOTA results in robustness.
Appendix A Ethics Statement
We exclusively employed publicly available datasets and models for research purposes and did not involve human subjects. EB-JDAT is a robust generative discriminator with powerful image-generation capabilities; if misused, it could produce content that poses societal risks, and no safeguards currently limit its outputs. Nevertheless, our approach achieves significantly greater robustness than earlier generative discriminators, ensuring its safe deployment in real-world applications. We therefore contend that the benefits of our work far outweigh any potential harms arising from disclosure of our methods. In addition, a fundamental classifier endowed with generative capacity and discriminative robustness more effectively mitigates data scarcity and adversarial perturbations.
Appendix B Experimental details
The experiments were carried out on a platform featuring an AMD EPYC 7542 32-core CPU with a clock speed of 2039.813 GHz and an NVIDIA GeForce RTX 3090 GPU, equipped with 24 GB of RAM. The proposed method was implemented using the open-source PyTorch machine learning framework. In addition, we use the same hyperparameter settings and validation-set configuration for training and evaluation as in JEM variants [2, 3].

### v2

@@ 4.3 Ablation Study
Ablation of Weighted Gradient Components.
In EB-JDAT, we employ a weighted combination of gradients, as detailed in Algorithm 1, to compute the overall optimization gradient, expressed as $h_{\theta}=w_{1}h_{1}+w_{2}h_{2}+w_{3}h_{3}$. To assess the individual contribution of each weight, we perform ablation studies by setting each of $w_{1}$, $w_{2}$, and $w_{3}$ to zero and evaluating their impact on overall performance, including (1) standard AT (NO.1) , (2) jointly modeling the energy distributions of adversarial and clean samples (NO.2) , and (3) jointly modeling the energy distributions of adversarial, clean, and generated samples (NO.3 and NO.4). The results, presented in Table 6, reveal two key findings: (1) $\frac{\partial\log p_{\theta}(\tilde{{\mathbf{x}}}|{\mathbf{x}})}{\partial\theta}$ ($h_{2}$) is not only the key to bridge the gap of accuracy, generation and robustness, but also stabilizes the training phase, reducing the risk of model collapse. (2) $\frac{\partial\log p_{\theta}({\mathbf{x}})}{\partial\theta}$ ($h_{1}$) can improve the ability of model in both classification and generation. Consequently, experimental results indicate that $w_{1}=1$, $w_{2}=1$, and $w_{3}=1$ provides the best trade-off between accuracy, robustness, generative capability, and stability. This configuration is therefore adopted as the default setting. Additionally, reducing $w_{2}$ can improve classification accuracy and generative quality, albeit with a slight decrease in robustness.
Ablation of Adversarial Sampling Steps.
Further, we ablate the number of adversarial sampling steps K $\in${1,5,10,20} while fixing the perturbation budget and step size. As shown in Fig. 5, a setting of 5 steps provided the best trade-off, allowing the model to converge stably. This choice was made to balance two factors: finding sufficiently challenging adversarial examples and avoiding model collapse, which is a known issue for energy-based models on complex data, as we note in our limitations section .
Table 6: Ablation study of EB-JDAT-JEM++, ECO denotes the epoch at which collapse occurs. Results in bold indicate the best.
@@
1 0 0 1 88.95 62.96 173.53 41
2 0 1 1 89.84 64.69 42.57 n/a
3 1 0.5 1 90.39 64.09 40.12 n/a
4 1 1 1 90.37 64.61 39.67 n/a
Figure 5: Ablation study of adversarial sampling steps. ECO=100 indicates no collapse. Time denotes the per-epoch cost.
5 Limitation And Discussion
Training EB-JDAT on complex, high-dimensional data remains challenging. This challenge is also encountered by JEMs,including JEM [13], JEM++ [39] and SADAJEM [41]. This instability arises from the typically sharp probability distribution of real data in high-dimensional space, which leads to inaccurate guidance for image sampling in regions with low data density. Although the approach we propose is a general and flexible optimization framework for all JEMs, considering training stability, we recommend training within faster and more stable JEM variants, such as JEM++ and SADAJEM. Nevertheless, our method significantly enhances the robustness of JEMs, surpassing SOTA AT [45], while incurring only a slight degradation in accuracy and generative performance, thereby achieving the best overall trade-off among robustness (68.76%) , accuracy (90.39%) , and generation (FID=27.42) .
@@ 6 Conclusion
In this work, we systematically investigate the underlying causes of the performance gap between discriminate classifiers, robust classifiers [25, 44] and JEM-based classifiers [13, 41] in terms of accuracy, generative capability, and adversarial robustness, and propose a general and flexible optimization framework, called EB-JDAT to eliminate the inherent trilemma. Our experiments validate the effectiveness of these techniques across various benchmarks, demonstrating superior performance in achieving the triple trade-off, with particularly impressive SOTA results in robustness.